Assuming the same stages as in Example 5.11, explain the potential pipeline hazards (if any) in...

  1. Assuming the same stages as in Example 5.11, explain the potential pipeline hazards (if any) in each of the following code segments.
    1. X=R2+Y; R4= R2+X

Time

Period

1

2

3

4

5

6

X=R2+Y

R4=R2+X

Answer:

  1. R1=R2+X; X =R3+Y; Z =R1+X

Time

Period

1

2

3

4

5

6

R1=R2+X

X = R3+Y

Z=R1+X
Answer #1

a. Yes there is pipeline data hazard in the instruction X=R2+Y; R4= R2+X and this hazard is Read After Write. The reason is first instruction is trying to perform Write operation over variable X and second instruction is trying to Read that updated value.

Hence second instruction has to wait for the first instruction to complete in this case and this is called Read After Write dependency. So second instruction will begin only after completion of first instruction.

Time

Period

1 2 3 4 5 6 7 8 9
X=R2+Y IF ID EX MEM WB
R4 = R2+ X IF ID EX MEM WB

b. In the set of instructions :-

R1=R2+X;

X =R3+Y;

Z =R1+X

There is Write After Read dependency between first and second instruction because first instruction is performing read over data X which is being written by second instruction. Then there is Read after Write dependency between second and third instruction due since second instruction is writing some value on data X which will be Read by third instruction.

Time

Period

1 2 3 4 5 6 7 8 9 10 11 12 13
R1=R2+X IF ID EX MEM WB
X =R3+Y IF ID EX MEM WB
Z =R1+X IF ID EX MEM WB

Please comment for any clarification.